OpenStudy (anonymous):

Solve the exponential equation. Express the solution set in terms of natural logarithms. 83x = 2.2

OpenStudy (anonymous):

FKKJWKEJI44 Please use proper notation it only takes a second and makes it easier for everyone trying to help you it is common curtesy

OpenStudy (anonymous):

*courtesy

OpenStudy (anonymous):

I will help you if you type it out using proper notation

OpenStudy (anonymous):

Solve the exponential equation. Express the solution set in terms of natural logarithms. 8^3x = 2.2

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

welcome :)

OpenStudy (anonymous):

take the log of both sides of the equation remember log_u(x) = u^(y) is the same as u^(y) = x thus if log_8(8) = 8^(y) so 8^(1) = 8 thus log_8(8) = 1 so take log_(8) on both sides of the equation log_8(8^(3x)) = log_8(2.2) 3xlog_8(8) = log_8(2.2) = 3x(1) = log_8(2.2) x = log_8(2.2)/3

OpenStudy (anonymous):

if you have any questions ask

OpenStudy (anonymous):

Remember that I simply used the rule log(x^(y)) = ylog(x)

OpenStudy (anonymous):

to bring down the exponent

OpenStudy (anonymous):

okay thanks

OpenStudy (anonymous):

if you dont care for an explanation you can use https://www.wolframalpha.com to find the answer